OpenStudy (anonymous):

Which of the following is a conjugate for −10 + 3i? −10 − 3i 3i − 10 3i + 10 10 − 3i

OpenStudy (anonymous):

@helpme1.2 @HelpBlahBlahBlah

OpenStudy (helpblahblahblah):

D

OpenStudy (anonymous):

Which of the following is the conjugate of 10 − 3i? 10 + 3i 10 − 3i 3 + 10i 3 − 10i @HelpBlahBlahBlah

OpenStudy (anonymous):

I do not believe you are suppose to Multiply across by a negative to get the conjugate

OpenStudy (anonymous):

z = a + ib Conjugate of z = a -ib

OpenStudy (anonymous):

so it is d correct

OpenStudy (helpblahblahblah):

one sec

OpenStudy (anonymous):

ok

OpenStudy (helpblahblahblah):

http://answers.yahoo.com/question/index?qid=20090406061353AA4vucH yup i think its still D

OpenStudy (anonymous):

-10 -3i